# Optimizing Estimators of Squared Calibration Errors in Classification

### v1

@@ 4.2 Kernel Ridge Regression
where $h\left(p,p^{\prime}\right)=\left\langle g,\left(\phi\otimes\phi\right)\left(p,p^{\prime}\right)\right\rangle_{\mathscr{H}\otimes\mathscr{H}}$. Then, a closed-form minimizer can be found, which results in the predictor
@@
It holds that $h_{\mathrm{ukkr}}=h_{\mathrm{kkr}}$ if the kernel used for $h_{\mathrm{kkr}}$ incorporates the regularisation constant $\lambda$ or when $\lambda=0$ (Stock et al., 2018).
In the next section, we perform top-label confidence and canonical calibration evaluations with the discussed and proposed estimators. Specifically, we use the proposed calibration estimation risk in Equation (20) for comparison and the proposed calibration-evaluation pipeline of Section 3.2.2 for calibration estimation.
@@ 5 Experiments
In this section, we demonstrate how to use our proposed risk framework in practice. We first run a simulation with known ground truth. Then, we evaluate the risk of the different calibration estimation function defined in Section 4 and the respective estimated calibration error across a variety of image classification datasets and models. We focus on top-label confidence, since it is the most prominent, and canonical calibration, which is the most general. The source code is publicly available at https://github.com/waiting-for-acceptance.
@@
Model LeNet-5 Densenet-40 ResNetWide-32 Resnet-110 Resnet-110 SD
Estimator
[…]
@@ 5.1 Simulation
We construct a simulation experiment with known ground truth to demonstrate how our proposed risk identifies the solution. For this, we draw i.i.d. samples $P_{1},\dots,P_{500}\sim\operatorname{Dir}\left(\alpha_{1},\dots,\alpha_{5}\right)$ from a Dirichlet distribution with concentration parameters $\alpha_{1}=\dots=\alpha_{5}=0.04$. These samples represent the (in practice unknown) ground truth probability vectors of a classification task with 500 instances and 5 classes. Then, we sample $Y_{i}\sim P_{i}$ for $i=1\dots 500$ as target labels. We set the hypothetical model predictions to $f\left(X_{i}\right)=\operatorname{softmax}\left(\frac{3}{10}\log P_{i}\right)$ for $i=1\dots 500$, which represent miscalibrated predictions. The concentration parameters were set such that the model would have an accuracy of $\approx 90\%$. We then define a calibration estimation function $h_{\operatorname{sim}}\left(p,p^{\prime}\right)\coloneqq\left\langle p-\operatorname{softmax}\left(\frac{10}{3}\theta\log p\right),p^{\prime}-\operatorname{softmax}\left(\frac{10}{3}\theta\log p^{\prime}\right)\right\rangle$, which has a single learnable parameter $\theta\in\mathbb{R}$ referred to as temperature. The estimation function was chosen such that it matches the ground truth $h_{\operatorname{sim}}=h^{*}$ if and only if $\theta=1$. In Figure 1, we plot the mean results with standard deviations of the empirical risk according to 100 repetitions of the experiment. As can be seen, the empirical risk correctly identifies the correct calibration estimation function with $\theta=1$.
@@ Results
In Figure 2 we depict the corresponding $\operatorname{TCE}_{2}$ estimations. As can be seen, only $\operatorname{TCE}_{2}^{\operatorname{kde}}$ is occasionally an outlier relative to the other estimators, which is expected based on the reported risk values of Table 1 and Table 2. Even though, we cannot spot a direct connection between all risk values and the estimated calibration values in Figure 2, the large risk of $\operatorname{TCE}_{2}^{\operatorname{kde}}$ is indicative of a worse estimation. However, it is not surprising that differences in the risk do not always translate to differences in the estimated values, since the loss does not measure in which direction a calibration estimator function gives wrong predictions.
@@ Appendix A Overview
Here, we first give additional experimental results in Appendix B. The missing proofs are located in Appendix C.
@@ Appendix B Extended Experiments
We use the implementation of the calibration estimator function $h_{\operatorname{kde}}$ given by the original authors (Popordanoska et al., 2022b). For a small fraction of inputs, this implementation returns NaN as prediction. We remove these instances from the risk and calibration estimation calculation of $h_{\operatorname{kde}}$, which has a neglectable effect.
@@
In the following, we discuss the risks and calibration estimations of some left-out cases from the main paper.
@@
Table 4: Validation set square root risk ^$\sqrt{\hat{\mathscr{R}}_{\mathrm{CE}}}\times 100$ of $\operatorname{TCE}_{2}$ for different ImageNet models. Lower is better. The estimator $\operatorname{TCE}_{2}^{\operatorname{kde}}$ performs worse than the other estimators for DenseNet-161 and ResNet-152. For Pnasnet-5, all estimators perform similar.
@@
Table 5: Validation set risk ^$\sqrt{\hat{\mathscr{R}}_{\mathrm{CE}}}\times 100$ of $\operatorname{CCE}_{2}$ for CIFAR10 models. Lower is better. All estimators perform fairly similar.
@@ C.2 Proof for Theorem 2
$\begin{split}&\mathbb{E}\left[f\left(P,P\right)\right]\\ &=\int_{\mathbb{R}^{d}}f\left(p,p\right)\mathrm{d}\mathbb{P}_{P}\left(p\right)\\ &=\int_{\operatorname{supp}\left(\mathbb{P}_{P}\right)}f\left(p,p\right)\mathrm{d}\mathbb{P}_{P}\left(p\right)\\ &=\int_{\operatorname{int}\operatorname{supp}\left(\mathbb{P}_{P}\right)}f\left(p,p\right)\mathrm{d}\mathbb{P}_{P}\left(p\right)+\int_{\operatorname{bd}\operatorname{supp}\left(\mathbb{P}_{P}\right)}f\left(p,p\right)\mathrm{d}\mathbb{P}_{P}\left(p\right)\\ &=\int_{\operatorname{int}\operatorname{supp}\left(\mathbb{P}_{P}\right)\setminus D}f\left(p,p\right)\mathrm{d}\mathbb{P}_{P}\left(p\right)+\int_{\operatorname{bd}\operatorname{supp}\left(\mathbb{P}_{P}\right)}f\left(p,p\right)\mathrm{d}\mathbb{P}_{P}\left(p\right).\end{split}$ (37)

### v2

@@ 4.2 Novel Calibration Estimators based on Kernel Ridge Regression
where $h\left(p,p^{\prime}\right)=\left\langle g,\left(\phi\otimes\phi\right)\left(p,p^{\prime}\right)\right\rangle_{\mathscr{H}\otimes\mathscr{H}}$. This objective is in essence the risk in Equation (20) plus a regularization constant. Then, a closed-form minimizer can be found, which results in the predictor
@@
It holds that $h_{\mathrm{ukkr}}=h_{\mathrm{kkr}}$ if the kernel used for $h_{\mathrm{kkr}}$ incorporates the regularisation constant $\lambda$ or when $\lambda=0$ (Stock et al., 2018). We define the calibration estimators based on kernel ridge regression for $\operatorname{CCE}_{2}$ as
$\operatorname{CCE}_{2}^{\mathrm{kkr}}\left(f\right)\coloneqq\sqrt{\frac{1}{n}\sum_{i=1}^{n}h_{\operatorname{kkr}}\left(f\left(X_{i}\right),f\left(X_{i}\right)\right)}$ (33)
and
$\operatorname{CCE}_{2}^{\mathrm{ukkr}}\left(f\right)\coloneqq\sqrt{\frac{1}{n}\sum_{i=1}^{n}h_{\operatorname{ukkr}}\left(f\left(X_{i}\right),f\left(X_{i}\right)\right)}.$ (34)
We also use analogous estimators $\operatorname{TCE}_{2}^{\mathrm{kkr}}$ and $\operatorname{TCE}_{2}^{\mathrm{ukkr}}$ for estimating $\operatorname{TCE}_{2}$.
We have now established a large pool of possible calibration estimation functions we can compare and optimize. In the next section, we perform top-label confidence and canonical calibration evaluations with the discussed and proposed estimators. Specifically, we use the proposed calibration estimation risk in Equation (20) for comparison and the proposed calibration-evaluation pipeline of Section 3.2.2 for calibration estimation.
@@ 5 Experiments
In this section, we demonstrate how to use our proposed risk framework in practice. We first run a simulation with known ground truth. Then, we evaluate the risk of the different calibration estimation functions defined in Section 4 and the respective estimated calibration error across a variety of image classification datasets and models. We focus on top-label confidence since it is the most prominent, and canonical calibration, which is the most general. The source code is publicly available at https://github.com/SebGGruber/Optimizing_Calibration_Estimators.
@@
Figure 3: Average runtime with error bars of the estimators in Figure 2(a) on a single CPU thread. Optimizing the hyperparameters has a substantial computational cost.
@@ 5.1 Simulation
We construct a simulation experiment with known ground truth to demonstrate how our proposed risk identifies the solution. For this, we draw i.i.d. samples $P_{1},\dots,P_{500}\sim\operatorname{Dir}\left(\alpha_{1},\dots,\alpha_{5}\right)$ from a Dirichlet distribution with concentration parameters $\alpha_{1}=\dots=\alpha_{5}=0.04$. These samples represent the (in practice unknown) ground truth probability vectors of a classification task with 500 instances and 5 classes. Then, we sample $Y_{i}\sim P_{i}$ for $i=1,\dots 500$ as target labels. We set the hypothetical model predictions to $f\left(X_{i}\right)=\operatorname{softmax}\left(\frac{3}{10}\log P_{i}\right)$ for $i=1,\dots 500$, which represent miscalibrated predictions. The concentration parameters were set such that the model would have an accuracy of $\approx 90\%$. We then define a calibration estimation function $h_{\operatorname{sim}}\left(p,p^{\prime}\right)\coloneqq\left\langle p-\operatorname{softmax}\left(\frac{10}{3}\theta\log p\right),p^{\prime}-\operatorname{softmax}\left(\frac{10}{3}\theta\log p^{\prime}\right)\right\rangle$, which has a single learnable parameter $\theta\in\mathbb{R}$ referred to as temperature. The estimation function was chosen such that it matches the ground truth $h_{\operatorname{sim}}=h^{*}$ if and only if $\theta=1$. In Figure 1, we plot the mean results with standard deviations of the empirical risk according to 100 repetitions of the experiment. As can be seen, the empirical risk correctly identifies the correct calibration estimation function with $\theta=1$.
@@ Results
In Figure 2 we depict the corresponding $\operatorname{TCE}_{2}$ estimations. As can be seen, only $\operatorname{TCE}_{2}^{\operatorname{kde}}$ is occasionally an outlier relative to the other estimators, which is expected based on the reported risk values of Table 1 and Table 2. Even though, we cannot spot a direct connection between all risk values and the estimated calibration values in Figure 2, the large risk of $\operatorname{TCE}_{2}^{\operatorname{kde}}$ is indicative of a worse estimation. However, it is not surprising that differences in the risk do not always translate to differences in the estimated values, since the loss does not measure in which direction a calibration estimator function gives wrong predictions. Figure 3 shows the wall-clock runtime of each estimator in Figure 2(a) averaged across the classifiers for CIFAR10. Since all estimators only use the classifier outputs, the classifiers have no systematic influence on the runtime. As can be seen, using our optimization pipeline adds substantial computational costs to the calibration estimation. This is analogous to the costs of hyperparameter optimization for conventional machine learning models (Bischl et al., 2023).
@@
Model LeNet-5 Densenet-40 ResNetWide-32 Resnet-110 Resnet-110 SD
Estimator
[…]
@@ 6 Conclusion
Acknowledgments
The authors are very thankful to Achim Hekler, who computed the logits for the VisionTransformer experiments in this work. Further, the authors would like to thank David Holzmüller, Eugène Berta, and Florian Buettner for fruitful discussions regarding this work.
The majority of this work was conducted during a research visit at Inria, which was partly funded by a “DKFZ Cancer Research Academy Short-term Research Visit Grant”.
@@ Appendix A Overview
Here, we first give additional experimental content in Appendix B. The missing proofs are located in Appendix C.
@@ Appendix B Extended Experiments
In the following, we discuss further experimental details and results. Specifically, we evaluate the calibration of VisionTransformer classifiers (Dosovitskiy et al., 2020) trained on MedMNIST datasets (Yang et al., 2021) according to our approach.
@@
We use the implementation of the calibration estimator function $h_{\operatorname{kde}}$ given by the original authors (Popordanoska et al., 2022b). For a small fraction of inputs, this implementation returns NaN as the prediction. We remove these instances from the risk and calibration estimation calculation of $h_{\operatorname{kde}}$, which has a neglectable effect.
@@
All experiments are run on an Intel(R) Xeon(R) Gold 5218R with 2.1 GHz and a Macbook Pro M1.
@@
In the following, we discuss the risks and calibration estimations of some cases left out of the main paper and VisionTransformer classifiers.
@@
We train the VisionTransformer architecture (Dosovitskiy et al., 2020) on the MedMNIST classification datasets Blood, OCT, and Derma (Yang et al., 2021). Specifically, we use a pre-trained classifier from Huggingface [^2] https://huggingface.co/timm/vit_base_patch16_224.orig_in21k_ft_in1k (Accessed on 2nd Jan 2025) and fine-tune with a modification of (Capelle, 2022). The respective risk values are in Table 6. The kernel density estimator performs worse than the other approaches. No definitive winner can be declared due to the wide error bounds. In Figure 6, we show the associated calibration estimates. Similar to before, the risk values do not link to the calibration estimates.
@@
Table 4: Validation set square root risk ^$\sqrt{\hat{\mathscr{R}}_{\mathrm{CE}}}\times 100$ of $\operatorname{TCE}_{2}$ for different ImageNet models. Lower is better. The estimator $\operatorname{TCE}_{2}^{\operatorname{kde}}$ performs worse than the other estimators for DenseNet-161 and ResNet-152. For Pnasnet-5, all estimators perform similarly.
@@
Table 5: Validation set risk ^$\sqrt{\hat{\mathscr{R}}_{\mathrm{CE}}}\times 100$ of $\operatorname{CCE}_{2}$ for CIFAR10 models. Lower is better. All estimators perform similarly.
@@
Table 6: Validation set risk ^$\sqrt{\hat{\mathscr{R}}_{\mathrm{CE}}}\times 100$ of $\operatorname{TCE}_{2}$ estimators for VisionTransformer models on various MedMNIST datasets. Lower is better. The error bounds are too wide to make a definitive statement about which estimator is best. Contrary, we can claim that the kernel density estimator performs worst on the Blood and Derma datasets.
Dataset Blood Derma OCT
Estimator
[…]
Figure 6: Different top-label confidence calibration estimates of VisionTransformer models trained on MedMNIST datasets. Most calibration estimates approximately agree with each other. Like in previous instances, this demonstrates that the calibration estimate is not directly linked to the respective risk value in Table 6.
@@ C.2 Proof for Theorem 2
$\begin{split}&\mathbb{E}_{P}\left[f\left(P,P\right)\right]\\ &=\int_{\mathbb{R}^{d}}f\left(p,p\right)\mathrm{d}\mathbb{P}_{P}\left(p\right)\\ &=\int_{\operatorname{supp}\left(\mathbb{P}_{P}\right)}f\left(p,p\right)\mathrm{d}\mathbb{P}_{P}\left(p\right)\\ &=\int_{\operatorname{int}\operatorname{supp}\left(\mathbb{P}_{P}\right)}f\left(p,p\right)\mathrm{d}\mathbb{P}_{P}\left(p\right)+\int_{\operatorname{bd}\operatorname{supp}\left(\mathbb{P}_{P}\right)}f\left(p,p\right)\mathrm{d}\mathbb{P}_{P}\left(p\right)\\ &=\int_{\operatorname{int}\operatorname{supp}\left(\mathbb{P}_{P}\right)\setminus D}f\left(p,p\right)\mathrm{d}\mathbb{P}_{P}\left(p\right)+\int_{\operatorname{bd}\operatorname{supp}\left(\mathbb{P}_{P}\right)}f\left(p,p\right)\mathrm{d}\mathbb{P}_{P}\left(p\right).\end{split}$ (39)